The freezing point of benzene, C6H6, is 5.50°C at 1 atmosphere. A nonvolatile, nonelectrolyte that dissolves in benzene is DDT . A student dissolves 14.18 grams of DDT, C14H9Cl5 (354.5 g/mol), in 159.0 grams of benzene. Use the table of boiling and freezing point constants to answer the questions below. Solvent Formula Kb (°C/m) Kf (°C/m) Water H2O 0.512 1.86 Ethanol CH3CH2OH 1.22 1.99 Chloroform CHCl3 3.67 Benzene C6H6 2.53 5.12 Diethyl ether CH3CH2OCH2CH3 2.02 The molality of the solution is __ m. The freezing point of the solution is__ °C.
Answer #1

1)

mass of DDT = 14.18 g

moles of DDT = mass / molar mass

                      = 14.18 / 354.5

                      = 0.04 mol

molality = moles / mass of solvent in kg

            = 0.040 / 0.159

molality = 0.252 m

2)

delta Tf = Kf x m

To - Tf = Kf x m

5.50 - Tf = 2.53 x 0.252

Tf = 4.86 oC

freezing point = 4.86 oC